Country A has a growth rate of 4.8​% per year. The population is currently 4,274​,000, and the land area of Country A is 20,000,000,000 square yards. Assuming this growth rate continues and is​ exponential, after how long will there be one person for every square yard of​ land?
I have no clue how to start this. Can anyone help?

Answer by stanbon(75887)

Country A has a growth rate of 4.8​% per year. The population is currently 4,274​,000, and the land area of Country A is 20,000,000,000 square yards. Assuming this growth rate continues and is​ exponential, after how long will there be one person for every square yard of​ land?
I have no clue how to start this. Can anyone help?
-----
P(t) = 4,274,000*(1.048)^t
-----
Solve:: 4,274,000*1.048^t = 20,000,000
---
1.048^t = 4.67946
---
t = log(4.67946)/log(1.048)
-----
t = 32.9 yrs
-----------
Cheers,
Stan H.
--------